### v1

# Benchmark Leakage Trap: Can We Trust LLM-based Recommendation?

Abstract
The expanding integration of Large Language Models (LLMs) into recommender systems poses critical challenges to evaluation reliability. This paper identifies and investigates a previously overlooked issue: benchmark data leakage in LLM-based recommendation. This phenomenon occurs when LLMs are exposed to and potentially memorize benchmark datasets during pre-training or fine-tuning, leading to artificially inflated performance metrics that fail to reflect true model performance. To validate this phenomenon, we simulate diverse data leakage scenarios by conducting continued pre-training of foundation models on strategically blended corpora, which include user-item interactions from both in-domain and out-of-domain sources. Our experiments reveal a dual-effect of data leakage: when the leaked data is domain-relevant, it induces substantial but spurious performance gains, misleadingly exaggerating the model’s capability. In contrast, domain-irrelevant leakage typically degrades recommendation accuracy, highlighting the complex and contingent nature of this contamination. Our findings reveal that data leakage acts as a critical, previously unaccounted-for factor in LLM-based recommendation, which could impact the true model performance. We release our code at https://github.com/yusba1/LLMRec-Data-Leakage.
Benchmark Leakage Trap: Can We Trust LLM-based Recommendation?
Mingqiao Zhang Nanjing University 231220059@smail.nju.edu.cn          Qiyao Pengthanks: Corresponding Author. Tianjin University qypeng@tju.edu.cn          Yumeng Wang Tianjin University ymwang@tju.edu.cn
Chunyuan Liu Lightwheel lcy370884579@gmail.com          Hongtao Liu Du Xiaoman Financial Technology htliu@tju.edu.cn
1 Introduction
Recently, the Large Language Models (LLMs) (e.g., GPT-4 OpenAI (2023), LLaMA Touvron et al. (2023)) demonstrates unprecedented capabilities in natural language understanding, which has been employed in various fields. The recommender systems based on large language models (LLMs) have emerged as a promising direction Wu et al. (2023a); Wang et al. (2024). These models are pre-trained on massive corpora and encode rich knowledge information, which could enhance the recommender systems by deeper understanding of item content, user intent and contextual nuances Geng et al. (2022b); Li et al. (2023a); Liu et al. (2023); Geng et al. (2022a). For example, BinLLM Zhang et al. (2024) integrates collaborative information into LLMs using a text-like encoding strategy for recommendation, which has demonstrated superior performance.
However, should the results of these LLM-based recommender systems really be considered reliable? Recent studies have revealed that large language models are susceptible to data leakage, where models inadvertently memorize and reproduce training data during inference Carlini et al. (2021); Hisamoto et al. (2020); Carlini et al. (2022). This phenomenon has been extensively documented in various domains, from text generation Carlini et al. (2023) to question answering Lee et al. (2022), where LLMs have been shown to reproduce verbatim passages from their training corpora. And it largely affects the evaluation of model performance Zhou et al. (2023). Although there is ample evidence that LLMs are exposed to benchmark data during pre-training and affect downstream inference resultsCarlini et al. (2021); Lee et al. (2022), it remains unclear how leaked LLMs affect downstream recommendation results when used as the backbone model of recommendation systems.
This raises our guess: integrating a memorization-prone LLM into the recommendation pipeline may cause the resulting LLM-based recommender systems to inherit and potentially amplify the data leakage characteristics inherent in the LLM. In fact, the core of recommender systems is user interest modeling and item characteristics learning Zhang et al. (2024); Rendle (2012). However, the inherent data leakage in LLMs can obscure the boundary between authentic user preferences and memorized data artifacts. This interference not only distorts the modeling of latent user interests, but also compromises the learning of item representations, as the model may rely on memorized associations rather than underlying and real characteristics. Such leakage could ultimately bias the evaluation of LLM-based recommender systems, as test results may reflect the model’s prior exposure to evaluation data rather than its true recommendation ability.
To investigate the impact of data leakage on LLM-based recommendation systems, we design a experimental framework that could simulate real-world leakage scenarios. First, we construct mixed leakage datasets with controlled proportions by blending in-domain data (sampled from the target evaluation datasets) with out-of-domain data (collected from six external sources). Second, to simulate leakage scenarios, we fine-tune the LLMs via Low-Rank Adaptation (LoRA) based on the blended leakage dataset, named as dirty LLM. This choice of LoRA provides a controlled proxy for investigating how benchmark knowledge injection biases recommendation performance while preserving the foundation model’s base capabilities. Finally, we evaluated performance by comparing downstream recommendation models built on clean and dirty LLMs in two categories. Specifically, we evaluate the impact of data leakage on (1) different model architectures, (2) various leakage compositions (pure in-domain, pure out-of-domain, and mixed), and (3) different types of out-of-domain data sources. In this way, we can analyze the factors that determine a model’s susceptibility or robustness to data leakage.
Our work makes the following key contributions:
  1. 1.
    We are the first to identify and empirically demonstrate the benchmark data leakage problem in LLM-based recommender systems, revealing how pre-exposed LLMs can compromise the integrity of downstream evaluation metrics.
  2. 2.
    We develop a novel methodology for simulating realistic data leakage scenarios through controlled fine-tuning on strategically composed mixed-domain datasets, enabling investigation of how different types and degrees of data exposure affect recommendation performance.
  3. 3.
    Through extensive experiments, we demonstrate that: (i) impacts vary substantially across different model architectures and their resilience to leakage, (ii) in-domain leakage creates deceptive performance gains that mask true model capability, while (iii) out-of-domain contamination typically impairs performance, revealing a nuanced dual-effect phenomenon that challenges current evaluation practices.
2 Related Work
2.1 Large Language Models
Large Language Models (LLMs) are a class of foundational models pre-trained on vast corpora of text data, demonstrating remarkable capabilities in understanding and generating human-like text OpenAI (2023); Bommasani et al. (2021); Brown et al. (2020). These models, typically based on the Transformer architecture, learn to predict the next token in a sequence, thereby capturing intricate patterns, facts, and reasoning abilities embedded in the training data Wei et al. (2022); Kaplan et al. (2020). A significant security and privacy concern associated with these models is data leakage, wherein LLMs can store and subsequently reproduce verbatim content from their training datasets Carlini et al. (2021). The study of data leakage in LLMs has become a critical area of research. Seminal work by Carlini et al.  Carlini et al. (2021) first systematically demonstrates that LLMs can memorize and regurgitate training data, showing that simple prompting strategies could recover exact sequences.  Zhou et al. (2023) shows that data leakage of large models will interfere with subsequent performance evaluation. And in response to address this problem, mitigation strategies like differential privacy and training data deduplication have been proposed to reduce these risks Lee et al. (2022).
2.2 LLM-based Recommendation
The field of LLM-based recommendation has evolved rapidly, with researchers exploring various architectural paradigms to integrate language models with recommendation tasks Wu et al. (2023b); Lin et al. (2023). LLMRec and LLMRec+Collab. are two primary architectural paradigms, which are directly represented in our experimental evaluation.
The first category, LLMRec methods, leverages the inherent language understanding capabilities of LLMs directly, with minimal architectural modification. This category encompasses a spectrum of techniques: ICL Dai et al. (2023) operates in a zero-shot manner through in-context learning alone, Prompt4NR Zhang and Wang (2023) explores generic soft-prompt tuning for the recommendation task, while TALLRec Bao et al. (2023) specializes the model via instruction tuning.
In contrast, the second category, LLMRec+Collab. methods, explicitly integrates collaborative filtering signals to enhance the LLM’s reasoning with user-item interaction patterns. The integration strategies, however, vary significantly: PersonPrompt Li et al. (2023b) injects collaborative information by introducing trainable token embeddings for users and items, effectively performing personalized soft-prompt tuning. CoLLM Zhang et al. (2025), in its two instantiations, maps pre-trained collaborative embeddings from Matrix Factorization (CoLLM-MF) or a Deep Interest Network (CoLLM-DIN) into the LLM’s latent space. A distinctly different approach is taken by BinLLM Zhang et al. (2024), which encodes collaborative signals into a text-like format of binary hash codes for direct inclusion in the model’s prompt. This clear taxonomic distinction provides the essential context for interpreting our experimental results on how these fundamentally different paradigms respond to data leakage scenarios.
3 Methodology
This section details our experimental framework for investigating data leakage in LLM-based recommender systems. We first provide an overview of the experimental pipeline, followed by the construction of leakage data, the contamination process through LoRA adaptation, and the formulation of clean and dirty downstream recommenders.
Figure 1: Overview of the Experimental Framework. The workflow illustrates the construction of a Mixed Leakage Corpus through strategic sampling of In-Domain (ID) and Out-Of-Domain (OOD) data. The Clean Path represents the baseline evaluation using a frozen Base LLM, while the Dirty Path simulates benchmark contamination via LoRA injection. By comparing the Clean and Dirty Recommenders, we identify the Triple Effect of Leakage: spurious performance gains, stability, or degradation.
3.1 Overview
The complete workflow of our study is illustrated in Figure 1. To explore the effect of data leakage on recommendation models, we begin by sampling a portion of the original dataset to serve as leaked data, which contains a pre-defined ratio of in-domain (ID) and out-of-domain (OOD) samples depending on whether the data originate from the same domain as the evaluation benchmark.
We start from an open-source large language model (Vicuna-7B) that has already been pre-trained on large-scale text corpora and denote it as the Clean LLM (i.e., without leakage). To emulate contamination, we fine-tune this Clean LLM on the mixed leakage dataset using Low-Rank Adaptation (LoRA) [8]. Through this process, the model learns leakage-related patterns that are encoded into the newly added low-rank parameters, producing a Dirty LLM. The base parameters remain frozen throughout training, ensuring that only the LoRA modules carry the injected information.
Subsequently, both Clean and Dirty LLMs are employed as backbone encoders in downstream recommendation systems, forming Clean Recommenders and Dirty Recommenders respectively. All hyperparameters, data partitions, and model configurations are kept identical between the two to ensure a fair comparison. The performance differences between the two systems therefore directly reflect the influence of data leakage.
During evaluation, Clean Recommenders serve as baseline models for standard performance testing, while Dirty Recommenders are used to observe the potential distortion of benchmark results under contaminated conditions. By comparing the performance metrics (AUC/UAUC) of the Clean and Dirty Recommenders, we identify the Triple Effect of Leakage: (1) Spurious Gains induced by domain-relevant leakage, (2) scenarios that Remain Stable, and (3) performance Degradation caused by domain-irrelevant contamination. This design allows us to analyze how different leakage types and proportions affect model reliability in a controlled and reproducible way.
3.2 Leakage Data Construction
To simulate realistic benchmark leakage scenarios, we construct a mixed leakage dataset that combines both in-domain and out-of-domain sources. Each dataset is split into training, validation, and test sets following a 7:2:1 ratio.
First, we randomly sample 10% from the target dataset as in-domain leakage:
$D_{\mathrm{ID}}\sim\mathrm{Sample}\!\left(D_{\text{target}},\,p=0.1\right).$ (1)
Here, $D_{\mathrm{ID}}$ denotes the in-domain leakage subset drawn from the target dataset $D_{\text{target}}$ at proportion p=0.1p{=}0.1. The operator $\mathrm{Sample}(\cdot)$ denotes uniform sampling according to the stated proportion. This step fixes the reference scale for subsequent OOD construction.
Then, we draw OOD data from six external datasets (Epinions, Last.fm, MIND, Amazon-Sports, Amazon-Beauty, Gowalla), with the total OOD size equal to $6\times|D_{\mathrm{ID}}|$ and allocated equally across sources:
$D_{\mathrm{OOD}}^{(i)}\sim\mathrm{Sample}\!\left(D_{\text{ext}}^{(i)},\,|D_{\mathrm{ID}}|\right),\quad i\in\{1,\dots,6\},$ (2)
$D_{\mathrm{OOD}}\;=\;\bigcup_{i=1}^{6}D_{\mathrm{OOD}}^{(i)},\qquad|D_{\mathrm{OOD}}|\;=\;6\,|D_{\mathrm{ID}}|.$ (3)
In (2), each $D_{\mathrm{OOD}}^{(i)}$ is sampled from the $i$-th external dataset $D_{\text{ext}}^{(i)}$ with the same target size $|D_{\mathrm{ID}}|$, ensuring that every source contributes equally. The union in (3) aggregates the six OOD portions into $D_{\mathrm{OOD}}$, yielding a total size of $6\,|D_{\mathrm{ID}}|$. Here $|\cdot|$ denotes cardinality.
Finally, we combine ID and OOD to obtain the mixed leakage set:
$\begin{split}D_{\text{leak}}&=D_{\mathrm{ID}}\cup D_{\mathrm{OOD}},\\ |D_{\text{leak}}|&=|D_{\mathrm{ID}}|+|D_{\mathrm{OOD}}|=7\,|D_{\mathrm{ID}}|.\end{split}$ (4)
Equation (4) defines the final leakage corpus $D_{\text{leak}}$ as the combination of the in-domain portion $D_{\mathrm{ID}}$ and the aggregated out-of-domain portion $D_{\mathrm{OOD}}$, whose size follows directly from the construction above, i.e., $|D_{\text{leak}}|=7\,|D_{\mathrm{ID}}|$.
3.3 Injecting leaked data
We start from an already pretrained base LLM with parameters $\theta_{0}$. This model is used directly as the Clean LLM and its parameters remain frozen throughout the whole study. To obtain the Dirty LLM, we optimize only low-rank LoRA adapters on the mixed leakage dataset $D_{\text{leak}}$ (constructed in  4), while keeping $\theta_{0}$ unchanged.
Loss definition
Let $L_{\text{LLM}}(\cdot)$ denote the standard negative log-likelihood (NLL) under next-token prediction:
$L_{\text{LLM}}(\theta,D)=-\sum_{(x,y)\in D}\log P(y\mid x;\theta),$ (5)
where each pair $(x,y)$ consists of an input sequence $x$ and target tokens $y$, and $P(y\mid x;\theta)$ is the model likelihood parameterized by $\theta$. This loss is adopted only to train LoRA adapters; it does not modify the frozen base $\theta_{0}$.
Contaminated adaptation (Dirty LLM)
We train LoRA adapters on $D_{\text{leak}}$ as follows:
$\Delta\theta^{\text{leak}}_{\text{LoRA}}=\arg\min_{\Delta\theta_{\text{LoRA}}}\;L_{\text{LLM}}\!\big(\theta_{0}\oplus\Delta\theta_{\text{LoRA}},\;D_{\text{leak}}\big),$ (6)
$\theta^{\text{dirty}}_{\text{LLM}}=\theta_{0}\oplus\Delta\theta^{\text{leak}}_{\text{LoRA}}.$ (7)
Here, $\oplus$ denotes the injection of LoRA adapters into the frozen base (i.e., composing the base weights with low-rank additive updates). [^1] We use the same tokenizer and decoding head as the base model; only LoRA parameters are trainable. The optimization in (6) finds the adapter parameters that best fit the leakage corpus, thereby encoding leakage-specific patterns into $\Delta\theta^{\text{leak}}_{\text{LoRA}}$. Equation (7) defines the Dirty LLM as the composition of the unchanged base $\theta_{0}$ and the learned leakage adapters.
LoRA parameterization and what is updated
For any affine weight $W$ in the frozen base, the LoRA update is
$W_{\text{new}}=W+\frac{\alpha}{r}\,BA,$ (8)
where $B\in\mathbb{R}^{d\times r}$ and $A\in\mathbb{R}^{r\times k}$ are trainable low-rank matrices (rank $r\ll\min(d,k)$), and $\alpha$ scales the update magnitude. The product $BA$ is a rank-$r$ correction that adds to the frozen weight $W$; hence only $A,B$ are optimized, while all base weights $W\subset\theta_{0}$ remain fixed. This factorization provides two benefits: (i) parameter efficiency (few trainable parameters) and (ii) isolation of contamination (leakage information is confined within $\Delta\theta_{\text{LoRA}}$, enabling a clean comparison against the base model).
In subsequent experiments, the Clean LLM is simply $\theta_{0}$ (no adapters), and the Dirty LLM is $\theta^{\text{dirty}}_{\text{LLM}}$ from (7).
Justification for LoRA as a controlled proxy
We implement the continued pre-training mentioned in our Abstract via LoRA-based adaptation to achieve a controlled simulation. A potential concern is the alignment between LoRA-based simulation and the deep-seated data leakage occurring in original pre-training. We argue that LoRA Hu et al. (2021) serves as the most appropriate scientific instrument for this study for three fundamental reasons:
  • Isolation of Variables (Mechanistic Essence): Unlike full-parameter pre-training which introduces global distribution shifts, LoRA allows us to freeze the foundation model’s base weights $\theta_{0}$Ṫhis ensures a strict isolation of variables, where the observed performance shifts can be purely attributed to the “memorized artifacts” injected via adapters, rather than catastrophic forgetting or general capability drift of the base LLM Hu et al. (2021).
  • Representativeness of Downstream Leakage: While pre-training leakage is a core concern, real-world LLM-based recommenders are increasingly built via instruction-tuning or domain-specific adaptation of open-source models Bao et al. (2023). LoRA accurately reflects this ”fine-tuning contamination” stage, which is often the final and most direct point of inadvertent benchmark exposure before model deployment.
  • Conservative Lower-bound Estimation: As LoRA updates represent a tiny fraction of total model parameters, it serves as a conservative proxy. If even a lightweight parameter update can induce the significant “dual-effect” (spurious gains in ID vs. degradation in OOD) observed in our results, it provides a compelling lower-bound estimate of the far more pervasive and deep-rooted risks posed by full-parameter pre-training leakage.
Dataset Domain Interaction Type Users Items Interactions
ML-1M Movies User History 6,040 3,952 1,000,209
Amazon-Book Books User History 22,507 24,642 1,791,692
Epinions Products User History 49,290 139,738 664,824
Last.fm Music Interest Tags 1,892 17,632 92,834
MIND News Interest Tags 1,000,000 100,000 24,155,470
Amazon-Sports Sports Product Attributes 35,598 18,357 296,337
Amazon-Beauty Beauty Product Attributes 22,363 12,101 198,502
Gowalla Locations Geographical 196,591 1,128,161 6,442,890
Table 1: Dataset statistics used in our experiments.
Dataset Method Category AUC UAUC
Baseline After leak $\mathbf{\Delta}$AUC Baseline After leak $\mathbf{\Delta}$UAUC
ML- 1M ICL LLMRec 0.5271 0.4123 $\bm{-21.8\%}$ 0.5206 0.4034 $\bm{-22.5\%}$
Prompt4NR 0.7021 0.7456 $6.2\%$ 0.6679 0.7123 $6.6\%$
TALLRec 0.7036 0.6234 $\bm{-11.4\%}$ 0.6751 0.5923 $\bm{-12.3\%}$
CoLLM-MF LLMRec +Collab. 0.7228 0.7806 $\bm{+8.0\%}$ 0.6782 0.7456 $\bm{+9.9\%}$
CoLLM-DIN 0.7184 0.7456 $3.8\%$ 0.6801 0.7123 $4.7\%$
PersonPrompt 0.7151 0.6789 $5.1\%$ 0.6487 0.6123 $5.6\%$
BinLLM 0.7379 0.7823 $6.0\%$ 0.6908 0.7345 $6.3\%$
Amazon- Book ICL LLMRec 0.4748 0.3456 $\bm{-27.2\%}$ 0.4795 0.3423 $\bm{-28.6\%}$
Prompt4NR 0.7168 0.6234 $\bm{-13.0\%}$ 0.5804 0.5023 $\bm{-13.5\%}$
TALLRec 0.7302 0.7823 $7.1\%$ 0.5892 0.6345 $7.7\%$
CoLLM-MF LLMRec +Collab. 0.8034 0.8923 $\bm{+11.1\%}$ 0.6147 0.6823 $\bm{+11.0\%}$
CoLLM-DIN 0.8186 0.7823 $4.4\%$ 0.6401 0.6123 $4.3\%$
PersonPrompt 0.7209 0.7823 $\bm{+8.5\%}$ 0.5876 0.6345 $\bm{+8.0\%}$
BinLLM 0.8207 0.7456 $\bm{-9.1\%}$ 0.6262 0.5723 $\bm{-8.6\%}$
[…]
4 Experiment
In this section, we first describe the overall experimental settings, including datasets, baselines, and implementation details. Then, we present and analyze the experimental results to evaluate the effectiveness of our proposed method.
4.1 Experiment settings
4.1.1 Datasets
For evaluating the impact of LLM data leakage on downstream recommendation models, we conduct experiments on two widely-used recommendation datasets:
  • ML-1M Harper and Konstan (2015):MovieLens-1M recommendation benchmark, a widely used dataset for movie recommendation containing 1000209 timestamped user–movie ratings (1–5 scale) from 6040 users on 3900 movies. In addition to user-item interactions, it provides side information such as movie titles, etc.
  • Amazon-Book Ni et al. (2019b): This dataset is derived from the "Books" category of the well-known Amazon Product Review dataset. It comprises book purchasing and rating records, user-book interactions within the e-commerce domain.
For simulating the data leakage, we employ six external datasets from different domains to serve as out-of-domain leakage data, including Epinions (2010) (SNAP), Last.fm Celma (2010), MIND Wu et al. (2020), Amazon-Sports Ni et al. (2019a), Amazon-Beauty Ni et al. (2019a) and Gowalla (2011) (SNAP). The detailed information of the above datasets is presented in Table 1.
These OOD datasets are strategically selected based on two criteria: (i) Semantic Divergence: they cover diverse domains ranging from news to geographical check-ins to test the model’s resistance to non-target semantic noise; (ii) Structural Heterogeneity: they encompass varied interaction types (e.g., tags, attributes) to observe how the memorization of disparate behavioral patterns interferes with target recommendation logic.
4.1.2 Baselines
In this paper, we evaluate the model performance on two categories of LLM-based recommendation models: LLMRec models (ICL, Prompt4NR, TALLRec) and LLMRec models integrated collaborative information (PersonPrompt, CoLLM, BinLLM).
  • ICL Dai et al. (2023): This is a LLMRec method based on the In-Context Learning ability of LLM. It directly queries the original LLM for recommendations using prompts.
  • Prompt4NR Zhang and Wang (2023): This method uses both fixed and soft prompts to utilize traditional Language Models (LM) for recommendation. We extend this method to the Vicuna-7B for a fair comparison.
  • TALLRec Bao et al. (2023): A state-of-the-art instruction-tuned LLMRec method that aligns LLMs to the recommendation task via task-specific instructions and preference formats, improving zero/low-shot recommendation.
  • PersonPrompt Li et al. (2023b): An LLMRec method that injects collaborative signals by introducing new tokens and embeddings for users/items. It can be viewed as personalized soft-prompt tuning on top of an LLM.
  • CoLLM Zhang et al. (2025): A state-of-the-art framework that maps collaborative representations into the LLM latent space to couple LLM reasoning with collaborative signals. We use two instantiations: CoLLM-MF, where collaborative embeddings come from matrix factorization (MF), and CoLLM-DIN, where collaborative embeddings are extracted by Deep Interest Network (DIN) Zhou et al. (2018).
  • BinLLM Zhang et al. (2024): A text-like encoding approach that represents collaborative information as binary hash codes. BinLLM converts collaborative embeddings into text strings (binary codes) and directly incorporates them into the prompt, enabling seamless integration with LLM’s text processing capabilities.
Implementation Details All experiments are conducted on cloud computing servers equipped with NVIDIA RTX 5090 GPUs, each providing 32 GB of memory to support large-scale training. We build our implementation upon PyTorch 2.0 and the Hugging Face Transformers 4.39.3 library for efficient model deployment and experiment management. The base model used in all experiments is Vicuna-7B, containing approximately 7 billion parameters, and it is fine-tuned using the LoRA technique with rank $r=8$, scaling factor $\alpha=16$, and dropout rate of $0.05$. The training configuration includes a learning rate initialized at $2\times 10^{-4}$ and decayed to $1\times 10^{-5}$ following a cosine schedule. Each GPU handles a batch size of $2$, with gradient accumulation over $64$ steps to maintain stability and memory efficiency. The Adam optimizer is employed with a weight decay of $0.01$ throughout training.
4.1.3 Evaluation Metrics
Regarding evaluation metrics, we employ AUC and UAUC Liu et al. (2021) to assess the performance of models before and after data leakage. AUC is the area under the ROC curve that quantifies the overall prediction accuracy, which evaluates the overall ranking quality. UAUC is derived by first computing the AUC individually for each user over the exposed items and then averaging these results across all users, which provides insights into user-level ranking quality. Their ranges are both from 0 to 1. The larger the value, the better the performance. We compute the performance change:
$\Delta_{m}=\frac{Perf_{m}(\theta_{m}^{dirty})-Perf_{m}(\theta_{m}^{clean})}{Perf_{m}(\theta_{m}^{clean})}\times 100\%\ ,$ (9)
where $Perf_{m}$ represents the performance metric (AUC or UAUC). A positive value indicates performance improvement after data leakage, while a negative value indicates performance degradation.
4.2 Results and Analysis
4.2.1 Overall Experimental Results
We present the performance comparison of seven recommendation methods on ML-1M and Amazon-Book datasets under mixed data leakage (in-domain 10% + out-of-domain 60%) in Table 2. Overall, data leakage can affect the evaluation of model performance to some extent. We have the following observations.
First, the impact of data leakage on model performance is uncertain. Compared to the baseline, the recommender’s performance on the two datasets sometimes improves and sometimes deteriorates. In the ML-1M dataset, TALLRec’s AUC drops from 0.7036 to 0.6234 after injecting mixed data. And CoLLM-MF’s AUC upgrades from 0.7228 to 0.7806. The same model can perform differently on different datasets. For example, BinLLM upgrades from 0.7379 to 0.7823 but it drops from 0.8207 to 0.7456. It is also worth noting that the trends and magnitudes of AUC and UAUC are basically consistent. This indicates that data leaks have a roughly consistent impact on both overall and user-level recommendation tasks. So in the subsequent analysis, we only refer to the changes in AUC.
Second, the impact of data leakage on model performance evaluation cannot be ignored. Recommenders trained on LLMs with data leaks may change their original rankings when undergoing benchmark testing. In the Amazon-book dataset, PersonPrompt’s performance is not as good as BinLLM. But after injecting the AT dataset, PersonPrompt is superior to B under this benchmark.
Finally, different categories of models have varying sensitivities to data leakage. Recommendation models of the "LLMRec" class (Uses a large language model directly for recommendation) are more susceptible to the effects of hybrid databases compared with recommendation models of "LLMRec+Collab." class (Augments an LLM with collaborative filtering signals). LLMs are strong at modeling textual semantics, whereas recommendation relies heavily on collaborative signals from user–item interactions. Injecting or fine-tuning LLMs with collaborative information aligns them with the behavior-prediction objective, leading to better personalization, ranking accuracy, and robustness. Therefore, on the baseline, LLMRec+Collab. models perform better. After injecting hybrid data, their performance change is smaller than that of pure LLMRec models. We speculate that this is because pure LLMRec methods lack alternative signals to offset contaminated knowledge, making them more vulnerable to degradation. By contrast, LLMRec+Collab. methods integrate collaborative information as an additional signal that provides redundancy, cross-signal validation and robustness enhancement, yielding a system that is less dependent on potentially leaked LLM parameters.
4.2.2 Ablation Study
We also randomly generate a few other mixed datasets to simulate possible data leaks. In the following experiment, to save space, we just represent the performance of TALLRec, CoLLM-MF and BinLLM under the baseline ML-1M. It should be noted that due to the inconsistent impact of different data on the mode, the following results may not logically correspond exactly with the data in Table  2, but this does not prevent us from drawing conclusions.
ML-1M (pure ID interference: 10% ID data)
Metric AUC UAUC
Method $\DeltaAUC After leak Δ\DeltaUAUC$
TALLRec 0.8795 $25.0\%$ 0.8474 $25.5\%$
CoLLM-MF 0.8234 $13.9\%$ 0.7723 $13.9\%$
BinLLM 0.8456 $14.6\%$ 0.7923 $14.7\%$
Table 3: AUC/UAUC after leak and relative change rate $\Delta$ on ML-1M under the “LLM stage 10% data leakage” setting. ID denotes the In-Domain.
ML-1M (pure OOD interference: 60% OOD data)
Metric AUC UAUC
Method $\DeltaAUC After leak Δ\DeltaUAUC$
TALLRec 0.5234 $25.6\%$ 0.4923 $27.1\%$
CoLLM-MF 0.6345 $12.2\%$ 0.5923 $12.7\%$
BinLLM 0.6234 $15.5\%$ 0.5823 $15.7\%$
Table 4: AUC/UAUC after leak and relative change rate $\Delta$ on ML-1M under the “pure OOD 60%” setting. OOD denotes the Out-of-Domain.
In Tables  3 and  4, we study purely data breaches by only injecting ID (In-Domain) data or OOD (Out-of-Domain), which are two idealized situations. It can be observed that after a data breach, the performance of models in Table  3 improves, while the performance of models in Table  4 declines. This indicates that data leakage within the domain can improve the accuracy of model predictions, whereas the opposite is true for out-of-domain data.
ML-1M (mixed leakage: 5% in-domain + 30% out-of-domain)
Metric AUC UAUC
Method $\DeltaAUC After leak Δ\DeltaUAUC$
TALLRec 0.6812 $3.2\%$ 0.6487 $3.9\%$
CoLLM-MF 0.7656 $5.9\%$ 0.7123 $5.0\%$
BinLLM 0.7534 $2.1\%$ 0.7045 $2.0\%$
Table 5: ML-1M under mixed leakage (5% in-domain + 30% out-of-domain): AUC/UAUC after leak and relative change rate $\Delta$.
Figure 2: AUC under different data injection methods
ML-1M (mixed leakage: 25% in-domain + 45% out-of-domain)
Metric AUC UAUC
Method $\DeltaAUC After leak Δ\DeltaUAUC$
TALLRec 0.7456 $6.0\%$ 0.7123 $5.5\%$
CoLLM-MF 0.8123 $12.4\%$ 0.7623 $12.4\%$
BinLLM 0.7823 $6.0\%$ 0.7345 $6.3\%$
Table 6: ML-1M under mixed leakage (25% in-domain + 45% out-of-domain): AUC/UAUC after leak and relative change rate $\Delta$.
In Table  5, we change the total amount of the hybrid database while we adjust the proportion of data in the internal and external domains in Table  6. The impact of the mixed data from the experiment in Table  5 on the model is smaller, with the variation rates of all three models not exceeding 6%. In Table  6, after increasing the proportion of in-domain data, the performance of downstream models trained by data-contaminated LLMs improved. It can be seen that data leakage of different degrees has varying effects on model performance.
In Fig. 2 and  3, we conduct a research on the impact of different data injection methods on model performance. We select three databases with different data construction methods as the ODD data:
  • Epinions: Built based on user history (2010) (SNAP) (the same as the LM-1M).
  • Last.fm: Built based on user interest tags Celma (2010).
  • Amazon-Sports: Built based on product attributes Ni et al. (2019a).
We find that the models’ performance after injecting Epinions data is not significantly different from before the injection while data from Last.fm and Amazon-Sports significantly reduces the accuracy of the model. This is because data from Epinions and baselines are in the same format. Although LLMs remember them, they don’t know the answers like they do with test data, but these data also fail to cause much interference. As a result, the impact of different types of out-of-domain data leaks on the model is also inconsistent.
Figure 3: UAUC under different data injection methods
5 Discussion
Our experimental findings reveal a critical vulnerability in LLM-based recommender systems evaluation: benchmark data leakage can significantly distort performance metrics. The observed performance inflation with in-domain data contamination challenges the validity of reported improvements, creating what we term the "leakage trap" in recommendation evaluation. This deceptive enhancement obscures true model generalization and compromises fair comparisons. The domain-dependent nature of leakage effects—where out-of-domain contamination typically degrades performance—further complicates the evaluation landscape. The varying resilience across architectures suggests that models relying more on collaborative signals may offer better protection against leakage effects.
At the same time, we plan to extend our investigation along several important dimensions. First, we intend to simulate data leakage on multiple other LLMs with varying scales and architectures to understand how model capacity and design choices affect leakage susceptibility. Second, we will explore recommendation models based on different LLM backbones to examine whether certain foundational models exhibit inherent resistance to contamination effects. Third, we plan to investigate temporal aspects of data leakage by examining how the recency of leaked data impacts model performance. Finally, we will study cross-domain leakage effects in more complex real-world settings where multiple domains interact simultaneously.
Several promising research directions emerge from our work. First, developing effective contamination detection methods for LLM pre-training corpora represents an urgent need, potentially leveraging techniques from dataset provenance and membership inference attacks. Second, designing leakage-robust evaluation metrics that can distinguish between genuine improvements and memorization artifacts would significantly advance the field. This could involve developing new testing protocols that explicitly control for potential contamination. Third, exploring architectural designs that inherently resist leakage effects, particularly through better integration of collaborative filtering signals and domain adaptation techniques, warrants deeper investigation. Additionally, studying the interaction between data leakage and other evaluation challenges like popularity bias and fairness represents another important direction.
From a practical perspective, establishing standardized protocols for reporting data provenance and conducting leakage audits could help build more trustworthy evaluation practices. This includes developing benchmark screening tools that can identify potential overlaps between training data and evaluation sets, as well as creating guidelines for transparent reporting of data curation processes. Furthermore, investigating the ethical implications of data leakage in production recommendation systems—particularly regarding user privacy and system fairness—constitutes another crucial research avenue.
6 Conclusion
This study provides empirical evidence of benchmark data leakage in LLM-based recommender systems and yields three key conclusions. First, in-domain data leakage causes substantial performance inflation that misrepresents true model capability. Second, leakage effects are domain-dependent, with out-of-domain contamination generally leading to performance degradation. Third, architectural choices significantly influence susceptibility, as hybrid models incorporating collaborative signals demonstrate greater resilience than pure LLM-based approaches. These findings underscore the critical need for more rigorous evaluation methodologies that account for data leakage risks, ensuring that reported advancements in LLM-based recommendation reflect genuine progress rather than evaluation artifacts.

### v2

@@
 1 Introduction
-Recently, the Large Language Models (LLMs) (e.g., GPT-4 OpenAI (2023), LLaMA Touvron et al.
+Recently, the Large Language Models (LLMs) (e.g., GPT-4 Achiam et al.
+(2023), LLaMA Touvron et al.
 (2023)) demonstrates unprecedented capabilities in natural language understanding, which has been employed in various fields.
 The recommender systems based on large language models (LLMs) have emerged as a promising direction Wu et al.
-(2023a); Wang et al.
+(2024); Wang et al.
 (2024).
 These models are pre-trained on massive corpora and encode rich knowledge information, which could enhance the recommender systems by deeper understanding of item content, user intent and contextual nuances Geng et al.
-(2022b); Li et al.
-(2023a); Liu et al.
-(2023); Geng et al.
-(2022a).
+(2022); Li et al.
+(2024); Liu et al.
+(2023).
 For example, BinLLM Zhang et al.
@@
 This phenomenon has been extensively documented in various domains, from text generation Carlini et al.
-(2023) to question answering Lee et al.
+(2023) to question answering Kandpal et al.
 (2022), where LLMs have been shown to reproduce verbatim passages from their training corpora.
@@
 Although there is ample evidence that LLMs are exposed to benchmark data during pre-training and affect downstream inference resultsCarlini et al.
-(2021); Lee et al.
+(2021); Kandpal et al.
 (2022), it remains unclear how leaked LLMs affect downstream recommendation results when used as the backbone model of recommendation systems.
@@
 2.1 Large Language Models
-Large Language Models (LLMs) are a class of foundational models pre-trained on vast corpora of text data, demonstrating remarkable capabilities in understanding and generating human-like text OpenAI (2023); Bommasani et al.
+Large Language Models (LLMs) are a class of foundational models pre-trained on vast corpora of text data, demonstrating remarkable capabilities in understanding and generating human-like text Achiam et al.
+(2023); Bommasani et al.
 (2021); Brown et al.
@@
 (2023) shows that data leakage of large models will interfere with subsequent performance evaluation.
-And in response to address this problem, mitigation strategies like differential privacy and training data deduplication have been proposed to reduce these risks Lee et al.
+And in response to address this problem, mitigation strategies like differential privacy and training data deduplication have been proposed to reduce these risks Kandpal et al.
 (2022).
@@
 The field of LLM-based recommendation has evolved rapidly, with researchers exploring various architectural paradigms to integrate language models with recommendation tasks Wu et al.
-(2023b); Lin et al.
+(2024); Lin et al.
 (2025).
@@
 methods, explicitly integrates collaborative filtering signals to enhance the LLM’s reasoning with user-item interaction patterns.
-The integration strategies, however, vary significantly: PersonPrompt Li et al.
-(2023b) injects collaborative information by introducing trainable token embeddings for users and items, effectively performing personalized soft-prompt tuning.
+The integration strategies, however, vary significantly: Personalized recommendation Lyu et al.
+(2024) injects collaborative information by introducing trainable token embeddings for users and items, effectively performing personalized soft-prompt tuning.
 CoLLM Zhang et al.
@@
   • Amazon-Book Ni et al.
-(2019b): This dataset is derived from the "Books" category of the well-known Amazon Product Review dataset.
+(2019): This dataset is derived from the "Books" category of the well-known Amazon Product Review dataset.
 It comprises book purchasing and rating records, user-book interactions within the e-commerce domain.
-For simulating the data leakage, we employ six external datasets from different domains to serve as out-of-domain leakage data, including Epinions (2010) (SNAP), Last.fm Celma (2010), MIND Wu et al.
+For simulating the data leakage, we employ six external datasets from different domains to serve as out-of-domain leakage data, including Epinions, Last.fm Celma Herrada and others (2009), MIND Wu et al.
 (2020), Amazon-Sports Ni et al.
-(2019a), Amazon-Beauty Ni et al.
-(2019a) and Gowalla (2011) (SNAP).
+(2019), Amazon-Beauty Ni et al.
+(2019) and Gowalla.
 The detailed information of the above datasets is presented in Table 1.
@@
 (2023): A state-of-the-art instruction-tuned LLMRec method that aligns LLMs to the recommendation task via task-specific instructions and preference formats, improving zero/low-shot recommendation.
-  • PersonPrompt Li et al.
-(2023b): An LLMRec method that injects collaborative signals by introducing new tokens and embeddings for users/items.
+  • PersonPrompt Lyu et al.
+(2024): A LLMRec method that injects collaborative signals by introducing new tokens and embeddings for users/items.
 It can be viewed as personalized soft-prompt tuning on top of an LLM.
@@
 In Fig. 2 and  3, we conduct a research on the impact of different data injection methods on model performance. We select three databases with different data construction methods as the ODD data:
-  • Epinions: Built based on user history (2010) (SNAP) (the same as the LM-1M).
-  • Last.fm: Built based on user interest tags Celma (2010).
-  • Amazon-Sports: Built based on product attributes Ni et al. (2019a).
+  • Epinions: Built based on user history (the same as the LM-1M).
+  • Last.fm: Built based on user interest tags Celma Herrada and others (2009).
+  • Amazon-Sports: Built based on product attributes Ni et al. (2019).
 We find that the models’ performance after injecting Epinions data is not significantly different from before the injection while data from Last.fm and Amazon-Sports significantly reduces the accuracy of the model.